OpenStudy (anonymous):

2x^2+18x+28/x^2+2x-35

OpenStudy (anonymous):

Try changing them into factors first before dividing

OpenStudy (anonymous):

\[\frac{2(x^2+9x+14)}{(x+7)(x-5)}\]

OpenStudy (anonymous):

\[\frac{2(x+2)(x+7)}{(x+7)(x-5)}\]

OpenStudy (anonymous):

By the way, math doesn't suck

OpenStudy (anonymous):

x+7 cancle\[\frac{2(x+2)}{x-5}\]

OpenStudy (anonymous):

@heromiles It does for some people.

OpenStudy (anonymous):

don't pay tension